Solve using multiplication principle
-3/4x=-5/6
It makes no sense to me!
Thank you

Answer by stanbon(75887)

Solve using multiplication principle
(-3/4)x=(-5/6)
---
Multiply both sides by -4/3 to get:
-----
x = (-4/3)(-5/6)
---
x = 20/18
---
x = 10/9
===============
Cheers,
Stan H.
============